OpenStudy (anonymous):

In a change jar, Tom has nickels and dimes. The number of nickels is two times the number of dimes and the total amount of money in the jar is $1.20. How many dimes are in the jar?

OpenStudy (texaschic101):

.05n + .10d = 1.20 (multiply by 100 to get rid of decimals) 5n + 10d = 120 n = 2d now sub 2d in for n 5(2d) + 10d = 120 10d + 10d = 120 20d = 120 d = 120/20 d = 6 n = 2d n = 2(6) n = 12 check.. 5n + 10d = 120 5(12) + 10(6) = 120 60 + 60 = 120 120 = 120 (correct) there is 12 nickels and 6 dimes

OpenStudy (anonymous):

Thanks! And could you stay on I will probably run into another word problem.

OpenStudy (texaschic101):

I will be on for a little while longer :)

OpenStudy (anonymous):

Okay so the quotient of five and seven numerically would be 5 divided by 7 or 7 divided by 5?